Flying against the jetstream, a jet travels 3700 km in 5 hours. Flying with the jetstream, the same jet travels 9280 in 8 hours. What is the speed of the jet in still air, and what is the speed of the jetstream?

Answer by checkley77(12844)

3700/5=s-j
9280/8=s+j
-----------------
3700/5+9280/8=2s
(8*3700+5*9280)/40=2s
40*2s=29,600+46,400
80s=76,000
s=76,000/80
s=950 km/h is the speed of the jet.
3700/5=950-j
740-950=-j
j=210 speed of the jet stream.
proof:
9280/8=950+210
1160=1160